Write the equation in its equivalent exponential form. 2 = log -49
Answer #1

We have given, it into 2 =log, 49 and we have to convert exponential form, So, If we have, x = logay then exponential form is
Answer #2

SOLUTION :


2 = log7 (49)


Raise both sides to the base 7 :

=> 7^2 = 7^(log7 (49)) 

=> 7^2 = 49 (ANSWER).


answered by: Tulsiram Garg